### v1

# October 2024

Abstract.  Addressing global societal challenges necessitates insights and expertise that transcend the boundaries of individual disciplines. In recent decades, interdisciplinary collaboration has been recognised as a vital driver of innovation and effective problem-solving, with the potential to profoundly influence policy and practice worldwide. However, quantitative evidence remains limited regarding how cross-disciplinary efforts contribute to societal challenges, as well as the evolving roles and relevance of specific disciplines in addressing these issues. To fill this gap, this study examines the long-term evolution of interdisciplinary contributions to the United Nations’ Sustainable Development Goals (SDGs), drawing on extensive bibliometric data from OpenAlex. By analysing publication and citation trends across 19 research fields from 1970 to 2022, we reveal how the relative presence of different disciplines in addressing particular SDGs has shifted over time. Our results also provide unique evidence of the increasing interconnection between fields since the 2000s, coinciding with the United Nations’ initiative to tackle global societal challenges through interdisciplinary efforts. These insights will benefit policymakers and practitioners as they reflect on past progress and plan for future action, particularly with the SDG target deadline approaching in the next five years.
Keywords.  Interdisciplinarity — Sustainability — Millennium Development Goals (MDGs) — Sustainable Development Goals (SDGs) — Agenda 2030 — Open Bibliometrics
{mdframed}
[linecolor=magenta] [Uncaptioned image] This version is a preprint of a paper authored by K. Okamura, titled ‘Evolving interdisciplinary contributions to global societal challenges: A 50-year overview’, posted on arXiv. The supplementary materials accompanying this manuscript have been seamlessly integrated into this single file.
1 Introduction: Interdisciplinary efforts for global societal challenges
It is widely recognised that the insights and expertise of a single academic discipline are insufficient to address global challenges effectively. Consequently, the importance of integrating knowledge and efforts across disciplines is increasingly acknowledged as essential to tackling these challenges. A substantial body of prior research has demonstrated the pivotal role of interdisciplinary collaboration in fostering innovation and addressing societal problems (Adler et al.,, 2018; Hu et al.,, 2024; Ledford,, 2015; Repko and Szostak,, 2020). Over the past few decades, numerous policy documents from national and international governmental organisations (Allmendinger et al.,, 2015; Gleed and Marchant,, 2016), alongside initiatives from academia (Aldrich,, 2014; NASEM,, 2005), and industry (Kotiranta et al.,, 2020; Suárez Giri and Sánchez Chaparro,, 2023), have drawn on these experiences and lessons to shape their objectives and strategies. Notably, at the Millennium Summit in September 2000, the United Nations (UN) Millennium Declaration led to the establishment of eight Millennium Development Goals (MDGs) (UN General Assembly,, 2000). This was followed in September 2015 by the adoption of 17 Sustainable Development Goals (SDGs) at the UN Summit on Sustainable Development (UN General Assembly,, 2015). Both frameworks are grounded in cross-disciplinary principles, motivations, and approaches.
In response to these global trends, the visibility and impact of publications related to interdisciplinarity have expanded markedly in recent years. Figure 1 illustrates the proportion of publications indexed in OpenAlex (Priem et al.,, 2022)—one of the largest catalogues of the world’s scholarly research—that contain the terms ‘multi(-)disciplinary’, ‘inter(-)disciplinary’, or ‘trans(-)disciplinary’ in their titles or abstracts, plotted by year of publication. The plots were generated for each of the four broad field categories, referred to as domains in OpenAlex. The graph highlights a noteworthy increase in the prevalence of these terms within the Social Sciences over the past few decades. The other three domains—Physical Sciences, Health Sciences, and Life Sciences—show a generally consistent upward trend, albeit with some variation in absolute numbers. This finding aligns with similar analyses conducted a decade ago (Larivière and Gingras,, 2014; Van Noorden,, 2015).
Figure 1: The number of publications related to ‘interdisciplinarity’ by publication year and domain. [Source: OpenAlex]
Nevertheless, there remains a lack of quantitative evidence on how interdisciplinary efforts have practically contributed to addressing global societal challenges. Several studies employing bibliometric approaches have analysed the relationship between societal challenges and various scientific subfields or relevant scientific journals, or have compared national contributions (Bedard-Vallee et al.,, 2023; Garai et al.,, 2023; Mishra et al.,, 2024; Raman et al.,, 2024; Sweileh,, 2020; Yamaguchi et al.,, 2023; Yumnam et al.,, 2024). These studies offer valuable insights into the connections between academia, the international community, and the SDGs. However, examining long-term contributions to global societal challenges from an interdisciplinary or cross-disciplinary perspective remains challenging, especially across the period encompassing both pre- and post-UN initiatives, including the MDGs and SDGs.
The knowledge and strategies that are effective in addressing specific societal challenges can differ significantly depending on the nature of the issue concerned. It is entirely possible for individuals from completely different disciplines to work on the same societal challenge without recognising each other’s efforts, applying knowledge and methods unique to their respective disciplines. For instance, which research fields have contributed to SDG 5, which aims at achieving gender equality and empowering all women and girls? Fields like sociology, political science, or psychology may immediately come to mind. However, as shown later (see Figure 3), medicine has made a significant contribution, especially in recent years—a fact that may not be readily apparent compared to the aforementioned fields. Each societal challenge often involves significant contributions from disciplines that may not initially be given much consideration. Understanding this reality and its temporal trends offers a vital perspective on interdisciplinary efforts and their contributions to various societal challenges. This understanding helps reveal the organic collaborative relationships and structures between academia and the policy sphere that work towards addressing common global problems. Such insights are invaluable for shaping future collaboration across disciplines, sectors, and borders.
In practice, conducting such analyses has proven challenging, primarily due to the difficulty of assessing whether—and to what extent—a given contribution has helped address, let alone resolve, a societal challenge (D’Este and Robinson-García,, 2023; Sandes-Guimarães et al.,, 2022). These evaluations are inherently part of a complex social and political process, and the methods, timeframes, and social consensus required make it difficult to discuss the results with any degree of certainty—scientifically, politically, or practically. Nevertheless, in today’s era of rapidly advancing Open Science and Data Science (Dong et al.,, 2017; Wittenburg,, 2021), and with the increasing availability of diverse scientometric data, it is now more feasible than ever to gain indirect insights through extensive data analysis. This is the approach we adopt in this paper.
2 Methods: Tracing interdisciplinary contributions to SDGs via bibliometrics
2.1 The OpenAlex data
The data used in the current study were retrieved from OpenAlex (Priem et al.,, 2022) between January and April 2024. OpenAlex gathers information on various types of publications, including journal articles, non-journal articles, reviews, preprints, conference papers, books, and datasets. This large-scale data catalogue employs the latest artificial intelligence (AI) technologies to analyse the disciplines closely related to various societal challenges and how these relationships have evolved over time. As a focal point for global societal challenges, we adopt the UN’s SDGs, introduced earlier, as a prominent and contemporary framework. Using bibliometric methods, we explore how different disciplines in the natural and social sciences have collectively addressed or related to these SDGs.
For disciplines, we use the 19 level-0 fields from the ‘concept’ taxonomy in OpenAlex. These fields include: Political Science, Philosophy, Economics, Business, Psychology, Mathematics, Medicine, Biology, Computer Science, Geology, Chemistry, Art, Sociology, Engineering, Geography, History, Materials Science, Physics, and Environmental Science, collectively referred to as $\mathscr{D}$ hereafter. [^1] From 2024 onwards, OpenAlex has introduced new classification categories to replace ‘concepts’, including 4,516 ‘topics’ (https://help.openalex.org/how-it-works/topics), 252 ‘subfields’, 26 ‘fields’, and 4 ‘domains’. For the analysis in the current paper, however, we chose to continue using ‘concepts’ because it offers an optimal level of granularity for our purposes—neither too detailed with too many fields nor too broad with too few fields—allowing for meaningful comparisons across a wide range of disciplines. For each of these fields, we collected the metadata for the top 1,000 most-cited publications from each publication year between 1970 and 2022. It is important to note that selecting publications based on citation counts does not imply that these publications are inherently superior in terms of scientific merit. Rather, we use citation counts as a proxy indicator of attention, reflecting the level of interest these publications have garnered from researchers and the wider community.
The number of publications examined in this study was $N=100,700$, derived from 1,000 publications per field per publication year, multiplied by 19 fields ($\mathscr{D}$) and 53 publication years (1970–2022). Refer to Supplementary Materials, including Tables S1 and S2, for further details on the OpenAlex data used in this study.
2.2 The field-affinity score
The metadata for each publication in OpenAlex includes a score indicating how closely the publication relates to the 19 level-0 concepts, [^2] https://docs.openalex.org/api-entities/works/work-object#concepts. which we refer to as the field-affinity score. These scores are calculated using specific deep learning algorithms (Barrett et al.,, 2023). The field-affinity score is assigned to each publication for each field as a number between 0 and 1, with 0 meaning the publication is unrelated to the field and 1 indicating maximal relevance to the field. Using the field-affinity scores, we calculated the distances between the 19 level-0 concepts ($\mathscr{D}$) for each publication year ($y$), following the same procedure outlined in Okamura, (2019), as follows.
To calculate the distance between fields A and B in $\mathscr{D}$, denoted as $d_{\mathrm{A}\mathrm{B}}$, we first identified and counted the number of publications ($n_{\mathrm{A}\cup\mathrm{B}}$) with a positive field-affinity score for either or both fields A and B from the total of $N=100,700$ publications. We then counted the number of publications ($n_{\mathrm{A}\cap\mathrm{B}}$) that had positive field-affinity scores for both fields A and B. Finally, we calculated the ratio of these counts and subtracted it from 1 to obtain the Jaccard distance between fields A and B:
$d_{\mathrm{A}\mathrm{B}}(y)=1-\frac{n_{\mathrm{A}\cap\mathrm{B}}(y)}{n_{\mathrm{A}\cup\mathrm{B}}(y)}\,.$ (1)
A value of 0 indicates that fields A and B are identical, while a value of 1 signifies that they are maximally distant.
Once the $19\times 19$ distance matrix for $\mathscr{D}$ is obtained, we can calculate the interdisciplinarity index for each publication. Specifically, labelling each publication by $i=1,\,\dots,\,N$ and denoting the set of publications from year $y$ as $S(y)$, the interdisciplinarity index for publication $i\in S(y)$ is calculated as follows:
$\mathit{\Delta}_{i}=\left(1-\sum_{\mathrm{A}\in\mathscr{D}}\sum_{\mathrm{B}\in\mathscr{D}}p_{i}^{\mathrm{A}}p_{i}^{\mathrm{B}}d_{\mathrm{A}\mathrm{B}}(y)\right)^{-1}\,.$ (2)
Here, we quantified interdisciplinarity using the so-called effective number of disciplines (Jost,, 2006; Leinster and Cobbold,, 2012; Okamura,, 2020). This index offers advantages over other commonly used diversity measures, such as the simple ‘variety’ index, Simpson’s index, or popular entropy measures like Shannon’s, in that it not only accounts for the number of fields and their composition (i.e., relative abundances) but also appropriately considers the distances between the constituent fields. Moreover, it is the most universal, as it can mathematically accommodate both broad categorisation and finer subdivisions of constituents in a consistent manner, satisfying the so-called Nesting Principle (Okamura,, 2020), which is not upheld by other measures such as the Rao-Stirling index (Rao,, 1982).
Using the interdisciplinarity index for each publication $i$, as defined by Eq. (2), the interdisciplinarity of field $\mathrm{A}\in\mathscr{D}$ is calculated as follows. First, for each publication year, the set of publications with a positive field-affinity score for field A is identified. Then, the mean of the $\mathit{\Delta}_{i}$ distribution, which is approximately bell-shaped, is calculated for field A within this set, and this value is taken as the interdisciplinarity index for field A:
$\mathit{\Delta}_{\mathrm{A}}(y)=\sum_{i\in S(y)}\mathit{\Delta}_{i}P(\mathit{\Delta}_{i})\,,$ (3)
where $P(\mathit{\Delta}_{i})$ denotes the probability of the interdisciplinarity variable taking the value $\mathit{\Delta}_{i}$ in the distribution. Using this index, we trace the evolution of each field’s interdisciplinarity, i.e., the degree to which it is associated with other fields.
2.3 The SDG-affinity score
[…]
To investigate the time evolution of the relative contribution of each field to each SDG, we first identify a set of publications, labelled by $i$, with positive affinity scores ($a_{i}>0$) for each SDG $m$ in each publication year ($y$). We then calculate the weighted sum of the affinity scores for each field ($\mathrm{A}\in\mathscr{D}$) across all publications belonging to this set, using the number of citations $(c_{i})$ for each publication as the weight, and denote this as $a_{\mathrm{A},m}(y)$. Specifically, denoting the set of publications with a positive ‘field A’-affinity score as $S(\mathrm{A})$ and those with a positive ‘SDG $m$’-affinity score as $S(m)$, we have:
$a_{\mathrm{A},m}(y)=\sum_{i\in S(y)\cap S(\mathrm{A})\cap S(m)}c_{i}a_{i}\,.$ (4)
In the above, the rationale for using the number of citations as a weight is based on the assumption that the attentional impact of a publication is proportionally associated with how many times it is subsequently cited. Then, the share of contribution from field $\mathrm{A}$ to SDG $m$ in year $y$ is evaluated as:
$R_{\mathrm{A},m}(y)=\frac{a_{\mathrm{A},m}(y)}{\sum_{m=1}^{M}a_{\mathrm{A},m}(y)}\,,$ (5)
satisfying $\sum_{m=1}^{M}R_{\mathrm{A},m}(y)=1$ for each year $y$. This index provides indirect insights into the relative increase (or decrease) in contributions from various fields, by publication year and by SDG.
3 Results and discussion: The shifting nexus of SDGs and disciplines
3.1 Disciplinary boundaries: Convergence with or isolation from other fields
Before examining the contributions of different fields to specific global societal challenges, we first present an analysis of how the degree of disciplinary convergence for each field, as measured by the index defined in Eq. (3), has evolved over time. The results of this analysis are shown in Figure 2. From the graph, we observe significant differences in the level of disciplinary convergence, as indicated by the interdisciplinarity index, across various fields. However, the most striking finding is not just the variance between fields, but the common trend seen across many: around the year 2000 (indicated by the dotted line), there is a distinct shift in the trajectory of the interdisciplinarity index for many fields.
Figure 2: Trends in interdisciplinarity across fields. For each field, a higher interdisciplinarity index score, calculated using Eq. (3), indicates a greater degree of interrelation with other fields. [Source: OpenAlex]
From around 1970 to 2000, although variations existed among fields, many showed a declining trend in interdisciplinarity. In fact, according to a simple linear regression analysis, the interdisciplinarity indices declined in 14 out of 19 fields ($p<0.001$). This may reflect a growing tendency for highly cited papers to become increasingly specialised within their respective disciplines, with mechanisms for generating new knowledge becoming more confined to these disciplines. In contrast, after around 2000, the declining trend in interdisciplinarity appeared to cease in many fields, and an overall increase in interdisciplinarity was observed across these fields. In fact, interdisciplinary indices rose in 10 out of 19 fields ($p<0.001$). The early 2000s coincide with the formulation of the MDGs at the UN and the initiation of related initiatives. The goals set at that time were difficult to achieve through the efforts of a single discipline alone and required collaboration across disciplines.
Assessing how directly this shift in interdisciplinarity indices reflects trends related to the global momentum associated with the MDGs and SDGs is complex. However, beyond these UN initiatives, Figure 2 strongly suggests that the global community, including academia, has increasingly introduced interdisciplinary perspectives, interests, or approaches into discussions. This phenomenon may also reflect the active publication of interdisciplinary research by high-impact journals and the proliferation of policy papers and other publications dealing with interdisciplinary issues. These observations are also consistent with the sharp increase in ‘interdisciplinarity’-related publications in Figure 1 since the 2000s.
3.2 Evolution of interdisciplinary contributions across SDGs
Building on the observed trend of disciplinary convergence across various fields, we now examine the interdisciplinary efforts directed towards each SDG and how the disciplinary composition has evolved over time through the analysis of the index defined in Eq. (5). The results are illustrated in Figure 3. This figure demonstrates that different global challenges are closely linked to knowledge from a wide range of fields. Furthermore, the degree of this linkage is unique to each challenge and has shifted dynamically over time. In some instances, the diversity of contributing fields has increased, while in others, the dominance of certain fields has grown, varying across different periods and disciplines.
Figure 3: Evolution of interdisciplinary contributions across the SDGs. For each SDG, the share of each field is calculated using Eq. (5). [Source: OpenAlex]
For instance, in the case of SDG 3 (Good Health and Well-Being), medicine has consistently accounted for over 50% of the contributions throughout the entire period, while biology has maintained approximately 20%. Although there is some degree of interdisciplinarity, this goal is predominantly influenced by these few key fields. This observation aligns with findings from an earlier study (Makarenko et al.,, 2021), which noted that interest in SDG 3 within academic circles is primarily focused on medical aspects, while economic dimensions are poorly represented. According to the cited literature, this relates to the underlying reasons for the lack of support and sufficient empirical evidence needed to address the financing challenges hindering the achievement of SDG 3.
In contrast, SDG 4 (Quality Education) has witnessed a significant increase in contributions from computer science over the past decade, despite its relatively steady presence in previous years. This trend may reflect ongoing debates regarding the impact of information and communication technology (ICT) and emerging digital technologies, such as Big Data, AI, and the Internet of Things (IoT), on sustainable education (Ghanem,, 2020; Tyagi et al.,, 2019). This surge in prominence for computer science is also observed in other SDGs, including SDG 9 (Industry, Innovation, and Infrastructure), SDG 10 (Reduced Inequalities), SDG 11 (Sustainable Cities and Communities), SDG 16 (Peace, Justice and Strong Institutions) and SDG 17 (Partnerships for the Goals). Given the rapidly evolving discussions surrounding the active use of AI, particularly in relation to achieving the SDGs (Gupta et al.,, 2021; Năstasă et al.,, 2024; Singh et al.,, 2024; UN ECOSOC,, 2024), it is likely that the contributions from information science to many SDGs—or their revised goals, if any, to be defined as successors—will continue to grow in the future, regardless of whether this is evident in graphs such as Figure 3.
For SDG 5 (Gender Equality), psychology accounted for approximately 40% of contributions until the mid-2000s, but this figure has since declined to around 20%. In contrast, the share of contributions from medicine has increased. While it is challenging to pinpoint the specific drivers behind this trend, discussions regarding the need to develop and strengthen health systems to support the women’s empowerment aspect of SDG 5 may have contributed. Furthermore, it is possible that medicine has partially merged with psychology to create a new interdisciplinary field relevant to SDG 5, which warrants further examination through expert knowledge. The consistent 10% contribution from political science throughout the period also reflects the political factors and interests surrounding this goal.
For SDG 7 (Affordable and Clean Energy), physics accounted for 30–40% of contributions before 2000, but its share has since declined to around 6%. In its place, materials science has gained prominence, accounting for approximately 40% of contributions since 2010. Chemistry has consistently contributed around 15%, while computer science has maintained a 15% share since 2000. This shift in the disciplinary composition may indicate a transition from a phase of scientific discovery to one more focused on engineering and practical applications in addressing energy issues. Other SDGs also offer distinct insights, each reflecting specific disciplinary characteristics as well as potential historical, social, cultural, industrial, or political contexts.
Comparing the panels in Figure 3, one observes that some SDGs exhibit an overall increasing trend towards disciplinary diversity, while others do not. To investigate this further, we conducted an additional analysis to examine the changes in the level of interdisciplinarity associated with each SDG. Specifically, for each SDG and each publication year, we identified publications with an SDG-affinity score above a certain threshold—here, a score of 0.5 was selected—and calculated the weighted mean interdisciplinarity index score for this set, using the number of citations for each publication as the weighting factor. The rationale for this approach remains consistent; in other words, the citation count serves as a proxy variable for attention. It is also important to note that this index abstracts the specific combinations and compositions of fields.
The results are displayed across all SDGs in Figure 4, revealing some intriguing trends characteristic of each SDG. For instance, as mentioned earlier, SDG 3, predominantly influenced by medicine, results in a consistently low interdisciplinarity index. In contrast, SDG 13 (Climate Action) has maintained a high level of disciplinary diversity, reflecting contributions from various fields, including environmental science, geography, geology, computer science, and physics. Meanwhile, SDG 9 (Industry, Innovation, and Infrastructure), which was influenced by fields such as mathematics and psychology in the 1970s, has recently seen a sharp rise in the prominence of computer science, leading to a marked decline in its interdisciplinarity index.
Figure 4: Trends in the Interdisciplinarity Index across the SDGs. For each SDG, the interdisciplinarity index is calculated based on its disciplinary composition, as depicted in Figure 3, with appropriate weighting, as well as the distances between fields, which are calculated using Eq. (1). [Source: OpenAlex]
4 Conclusions: Towards evidence-based actions for sustainable development
This paper has provided an overview of how various disciplines are connected to different global societal challenges and how these connections have evolved over more than half a century—an area that has previously lacked substantial quantitative evidence. Specifically, we utilised two types of scores based on specific machine learning algorithms employed in the OpenAlex platform: the field-affinity scores and the SDG-affinity scores. These scores were used to develop appropriate measures for analysing the organic and dynamic relationships between different disciplines and the SDGs.
Our analysis suggests that since the 2000s, when the MDGs were implemented in earnest, there has been a clear trend towards an increase in publications and/or citations with interdisciplinary content, with many fields showing a tendency to become more interconnected. The focus on specific global societal issues, as represented by the SDGs, has also highlighted that the interdisciplinarity of these issues and the ways in which they have changed over time are markedly different, reflecting the unique characteristics of each issue. A considerable body of literature has concentrated, and will continue to concentrate, on the analysis and discussion of each of the SDGs, particularly as we approach the five-year mark before the SDG target deadline. Within these ongoing discussions, the findings and visualisations presented in this study offer valuable and insightful perspectives from an interdisciplinary and cross-sectoral standpoint. These insights are especially relevant to policymakers and practitioners as they reflect on the past decades and strategise for the future.
Finally, we would like to address some considerations regarding the methodology employed in this study and outline expectations for future research. It is essential to continue evaluating and refining the approach used in this study, which applies an Open Bibliometrics-based lens to observe the complex landscape of diverse global societal challenges. Open data catalogues of scholarly publication data, including OpenAlex, utilised in this paper, are continually improving in both the quality and quantity of their collections, and these resources will increasingly serve as valuable tools for analysing and visualising global trends. While commercial databases like Web of Science and Scopus are beneficial for conducting precise analyses from specific perspectives, such open and freely accessible resources (Delgado-Quirós and Ortega,, 2024) can also offer valuable insights into broad trends, particularly for policy-related purposes (Kitajima and Okamura,, 2023; OECD.AI, n.d., ; Okamura,, 2023, 2024; UNESCO,, 2023). In these contexts, the emphasis is on quick comprehension and a general, albeit rough, understanding. By recognising the unique characteristics—including both strengths and limitations—of each scholarly database and by using or combining them appropriately for different applications, we can gain new insights and develop more effective lenses through which to examine global trends (Culbert et al.,, 2024; Haupka et al.,, 2024; Liu et al.,, 2023; Visser et al.,, 2021). Doing so will further enhance the quality and dynamism of discussions surrounding societal impact in both policy and academic spheres, contributing meaningfully to these critical dialogues.
Acknowledgements.
The views and conclusions contained herein are those of the author and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of any of the organisations with which the author is currently or has been affiliated in the past.
Competing Interests.
The author has no competing interests.
Funding Information.
The author did not receive any funding for this research.
Data Availability.
The datasets and figures generated and/or analysed during this study are available in the Zenodo repository at https://doi.org/10.5281/zenodo.13998979.
Supplementary Materials
The data derived from OpenAlex and used in this study, as outlined in the Methods section of the main text, are available in the Zenodo repository at https://doi.org/10.5281/zenodo.13998979. Figure 1 is based on the data presented in Table S1, while Figures 2, 3, and 4 are derived from the data in Table S2.
[…]
 Variable Description
 pyear Year of publication
 nwork Number of publications, indexed in OpenAlex
[…]
 nIDR Number of publications with ‘multi-/inter-/trans-disciplinar*’ in Title or Abstract, indexed in OpenAlex
[…]
 %nIDR Ratio of nIDR to nwork (in %)
[…]
 Variable Description
 idwork Work identifier starting with ‘W’, as defined in OpenAlex
 pyear Year of publication
 citation Number of citations, as of April 2024, indexed in OpenAlex
[…]
 SDGm Score of affinity to each of the United Nations’ Sustainable Development Goals (SDGs) ($\texttt{{\color[rgb]{1,0,0}m}}=\texttt{1},\,\dots,\,\texttt{17}$)